# Enhancing Few-Shot Image Classification through Learnable Multi-Scale Embedding and Attention Mechanisms

### v1

@@ Abstract
In the context of few-shot classification, the goal is to train a classifier using a limited number of samples while maintaining satisfactory performance. However, traditional metric-based methods exhibit certain limitations in achieving this objective. These methods typically rely on a single distance value between the query feature and support feature, thereby overlooking the contribution of shallow features. To overcome this challenge, we propose a novel approach in this paper. Our approach involves utilizing multi-output embedding network that maps samples into distinct feature spaces. The proposed method extract feature vectors at different stages, enabling the model to capture both global and abstract features. By utilizing these diverse feature spaces, our model enhances its performance. Moreover, employing a self-attention mechanism improves the refinement of features at each stage, leading to even more robust representations and improved overall performance. Furthermore, assigning learnable weights to each stage significantly improved performance and results. We conducted comprehensive evaluations on the MiniImageNet and FC100 datasets, specifically in the 5-way 1-shot and 5-way 5-shot scenarios. Additionally, we performed a cross-domain task from MiniImageNet to the CUB dataset, achieving high accuracy in the testing domain. These evaluations demonstrate the efficacy of our proposed method in comparison to state-of-the-art approaches. https://github.com/FatemehAskari/MSENet
@@ 1 Introduction
In recent years, deep learning has led to remarkable progress in the image recognition field, surpassing traditional computer vision algorithms bib1 ; fateh2024advancing . However, the success of deep learning models heavily relies on huge amount of data. When the available data is insufficient, the model faces difficulties in effectively optimizing its parameters bib2 , which can lead to overfitting. Consequently, achieving effective recognition becomes challenging, hindering the desired outcome rezvani2024single ; bib4 . Depending on the contexts, the process of labeling images can be time-consuming and costly fateh2024enhancing ; fateh2023persian . For this reason, having models that can achieve acceptable performance with a small number of samples is perceived to be of great importance bib7 .
@@
Data augmentation is one of the techniques used in situation where the amount of labeled data is insufficient bib8 . However, its impact on the model’s performance is limited because methods such as rotation, noise addition and scaling cannot provide new information and only help to some extent in preventing overfitting bib9 . Another technique is transfer learning bib10 , which transfer knowledge from a source domain to a target by freezing the shallow network and fine-tuning the deep network layers. However, if the target domain differs significantly from the source domain, this approach may not be very effective when dealing with a small number of samples bib11 .
To address the limitation of traditional deep learning models, which performed well with large training datasets, models were developed to perform well with a fewer number of samples. One of the most important approaches is meta-learning bib12 . Meta-learning aims to create models that can generalize across different tasks and rapidly adapt to new problem domain by leveraging prior learning experiences bib13 ; bib14 . The three main meta-learning approaches are: model-based meta-learning, optimization-based meta-learning, and metric-based meta-learning bib15 . Model-based methods, aiming for rapid learning, primarily concentrate on model architectures and adjust model parameters based on the given tasks. Commonly used architectures in model-based methods include convolutional neural networks (CNNs) and recurrent neural networks (RNNs) bib16 . Optimization-based methods aim to improve model initialization and gradient descent direction through episodic training, enabling effective learning with a small number of training samples. These methods consist of a task-specific learner and a meta-learner, trained on task distributions, to optimize model parameters for few-shot learning tasks bib17 ; bib18 . Metric-based methods focus on learning a distance metric that effectively measures the similarity between samples, with the goal of optimizing it for new learning tasks. In the context of few-shot image classification, these methods aim to ensure that samples from the same class have a small distance, while samples from different classes have a large distance bib19 ; bib20 .
Metric-based methods extract feature vectors and measure the distance between classes based on these extracted vectors. However, metric-based methods usually learn only one embedding space. Gao et al. bib21 proposed a multi-distance metric network that embeds samples into different spaces by extracting feature vectors at different stages. The shallow network layers capture global information about the images, while the deeper layers focus more on abstract features. Utilizing these different spaces has improved the performance of the model. In addition on different embedding spaces, it is also important to have feature vectors that are rich information. Afrasiyabi et al. bib22 proposed a set-feature extractor (SetFeat) that utilizes a self-attention mechanism to obtain a richer feature space.
@@
In this study, the proposed model integrates various components to achieve enhanced performance: The proposed method utilizes ResNet18 as the feature extractor, extracting feature maps from five different stages of ResNet18 to facilitate multi-scale representation. At each stage, we introduce learnable parameter weights to augment the model’s capacity for feature representation. Moreover, to enrich the feature space and further enhance multi-scale representation, self-attention mechanisms are incorporated into each feature map.
Though a comprehensive evaluation, we demonstrate the effectiveness of our model under MiniImageNet, CUB, and FC100 datasets.
Our contribution can be summarized as follow:
@@
    We extracted five feature maps from backbone in order to capture both global and task specific features.
@@
    We employ a self-attention mechanism for each feature map obtained from every stage in order to capture more valuable information.
@@
    We incorporate learnable weights at each stage.
@@
    We propose a novel few-shot classification. We have significantly improved the accuracy on the MiniImageNet and FC100 datasets.
@@ 2 Related Works
Koch et al. bib29 proposed a Siamese network that utilizes the VGG network as an extractor. They feed two pairs of images into the shared-weight convolutional network, and the network outputs a numerical value between 0 and 1, representing the similarity between the two images. Vinyals et al. bib24 proposed a matching network that computes the probability distribution over labels using an attention kernel. The attention kernel calculates the cosine similarity between the support set of embedded vectors and the query. It then normalizes the similarity using the softmax formula. Snell et al. bib30 proposed the Prototypical network, where each class in the support set is represented by a prototype, defined as the mean of the embedded vectors belonging to that class. The similarity between the query image’s embedded vector and the prototypes of each class is determined using the Euclidean distance. This enables the classification of query images into their respective classes. Sung et al. bib31 proposed the Relation Network, which does not rely on a separate distance function. Instead, it connects the representations of the support set and the query directly within the neural network architecture, allowing the network to learn the similarity measure. Previous few-shot image classification methods commonly used four-layer convolutional networks as backbones. However utilization of pre-trained networks such as ResNet12 and ResNet18 has become much more popular nowdays. However, calculating similarities and differences to a single feature vector is not sufficient. Gao et al. bib21 proposed a model called MDM-Net for few-shot learning. The MDM-Net maps input samples into four different feature spaces using a multi-output embedding network. Additionally, they introduced a task-adaptive margin to adjust the distance between different sample pairs. While previous approaches commonly used CNNs as feature extractors, there is a growing trend towards utilizing Transformers and attention mechanisms. CNNs have limitations such as a limited receptive field and parameter inefficiency. On the other hand, Transformers offer advantages including the ability to capture long-range dependencies, model non-local relationships, and parallelize computations efficiently. They also provide interpretability by highlighting important regions or features in the input data, aiding in the understanding of the model’s decision-making process. The MDM-Net introduces a promising approach in few-shot learning by leveraging the strengths of Transformers and attention mechanisms while addressing the limitations of CNNs.
@@
Wang et al. bib32 propose a unified Query-Support Transformer (QSFormer) model for few-shot learning. The QSFormer model addresses the challenges of consistent image representations in both support and query sets, as well as effective metric learning between these sets. It consists of a sampleFormer branch that captures sample relationships and conducts metric learning using Transformer encoders, decoders, and cross-attention mechanisms. Additionally, a local patch Transformer (patchFormer) module is incorporated to extract structural representations from local image patches. The proposed model also introduces a Cross-scale Interactive Feature Extractor (CIFE) as an effective backbone module for extracting and fusing multi-scale CNN features. The QSFormer model demonstrates superior performance compared to existing methods in few-shot learning. Ran et al. bib32 propose a novel deep transformer and few-shot learning (DT-FSL) framework for hyperspectral image classification. The framework aims to achieve fine-grained classification using only a few-shot instances. By incorporating spatial attention and spectral query modules, the framework captures the relationships between non-local spatial samples and reduces class uncertainty. The network is trained using episodes and task-based learning strategies to enhance its modeling capability. Additionally, domain adaptation techniques are employed to reduce inter-domain distribution variation and achieve distribution alignment.
@@ 3.1 Problem definition
Figure 1: SA module
The goal of few-shot classification is to classify a unseen sample. We have two datasets, $D_{train}$ and $D_{test}$, each associated with corresponding class sets $C_{train}$ and $C_{test}$, respectively. It is important the class sets $C_{train}$ and $C_{test}$ are disjoint, meaning they have no elements in common. Formally, we can express this as $C_{train}\cap C_{test}=\emptyset$. Each training episode consists of a support set $S$ and a query set $Q$. The support set $S$ comprises $K$ examples for each of $N$ distinct classes, denoted as $X_{i}^{S}=\left\{\left(x_{ij}^{s},C=i\right)\right\}_{i=1}^{K}$, where $x_{i_{j}}^{s}$ represents the $j$th example belonging to class $i$. The query set $Q$ contains $X^{q}=\left\{x_{i}^{q}\right\}_{i=1}^{n_{q}}$. The objective of the model is to leverage the support set $S$ to correctly classify the query example $x_{i}^{q}$. In other words, the model is trained to predict the class label of the query image based on the set of supporting examples and their class affiliations provided in $S$. During training, each episode consists of a random sample drawn from the training dataset $D_{train}$. The objective of the model is to learn to extract feature representations from the examples in the support set $S$, such that the distance between the feature vector of the query image $x_{i}^{q}$ and the feature vectors of the support examples $x_{i_{j}}^{s}$ can be effectively measured. Specifically, the model learns to extract discriminative feature vectors from the support examples, which can then be used to classify the query image based on its proximity to the support set features. This training paradigm encourages the model to rapidly adapt its feature extraction and classification capabilities from the limited support data to accurately predict the class label of the query instance. Similarly, during evaluation, the performance of the trained model is assessed on the held-out test dataset $D_{test}$. In this phase, the model extracts feature vectors for each example in the test set, leveraging the knowledge and feature extraction capabilities it learned during the training episodes on the $D_{train}$ dataset.
@@ 3.2 Overview
The overall architecture of our model is illustrated in fig. 2. Our proposed approach incorporates several key components designed to enhance performance. At the core is a robust backbone architecture that enables the extraction of feature maps across diverse spatial scales. Additionally, we have integrated an attention module to further refine the feature extraction process. Underpinning our framework is a distance metric that facilitates effective similarity computation between inputs. Moreover, we have incorporated learnable weights to capture the relative significance of each extracted feature map.
Figure 2: Final model
@@ 3.2.1 Backbone
We utilized a pre-trained ResNet-18 network with initial weights from the ImageNet dataset. We removed the last fully-connected layer and fine-tuned the network on our specific dataset. To leverage multi-level feature maps, we proposed a multi-output embedding approach where we extracted feature maps at the end of each of the 5 convolutional blocks in the ResNet-18 architecture. This allowed us to capture feature representations at multiple scales and resolutions. Each sample in the support set, denoted as $x_{ij}^{s}$, is mapped to five different feature spaces like $f_{ij}^{s}$ as shown in equation 1:
@@
In our proposed approach, we employ multiple convolutional layers $(Conv1,Conv2,Conv3,Conv4,Conv5)$ from the ResNet-18 network to extract features from the samples. Each of these convolutional blocks contributes to a distinct feature representation, allowing us to leverage different levels of abstraction and information contained in the feature maps for both the support and query samples. This multi-level feature extraction strategy enables a more comprehensive and robust representation of the input images, which is crucial for effective similarity computation and classification tasks.
@@ 3.2.2 Attention module
After extracting the feature vectors at each stage, we utilize self-attention and global average pooling.The representation of the attention module is illustrated in fig. 1. Considering the extracted feature vectors, we apply a $1\times 1$ convolution on $f$, resulting in convolutional vectors $k$, $g$ and $h$. This operation is performed to reduce the number of channels (equation 3):
@@
The attention mechanism computes weights $\beta$, that determine the relative importance of each pixel in the feature map. These weights are calculated across the entire spatial extent, allowing the attention module to capture long-range dependencies beyond a local neighborhood, unlike traditional convolutions. The output of the self-attention layer is computed as equation 5:
$o^{Conv-p}=Conv_{1\times 1}(\sum_{i=1}^{N}\beta_{i,j}v^{\prime}(f^{Conv-p}))$ (5)
In equation 5, in addition to using various ResNet18 blocks for feature extraction, we employ a $1\times 1$ convolution layer twice - once before applying the attention mechanism and again after it. The purpose of using the $1\times 1$ convolution layer twice is to make the network more nonlinear and to allow us to leverage the previous information more evenly. The scaling factor, typically denoted as $\gamma$, is a learnable parameter in the network that is multiplied with the input feature map $\mathbf{f}^{Conv-p}$ before the addition (equation 6):
$y^{Conv-p}=\gamma\times o_{i}+f^{Conv-p}$ (6)

### v2

@@ Abstract
In the context of few-shot classification, the goal is to train a classifier using a limited number of samples while maintaining satisfactory performance. However, traditional metric-based methods exhibit certain limitations in achieving this objective. These methods typically rely on a single distance value between the query feature and support feature, thereby overlooking the contribution of shallow features. To overcome this challenge, we propose a novel approach in this paper. Our approach involves utilizing a multi-output embedding network that maps samples into distinct feature spaces. The proposed method extracts feature vectors at different stages, enabling the model to capture both global and abstract features. By utilizing these diverse feature spaces, our model enhances its performance. Moreover, employing a self-attention mechanism improves the refinement of features at each stage, leading to even more robust representations and improved overall performance. Furthermore, assigning learnable weights to each stage significantly improved performance and results. We conducted comprehensive evaluations on the MiniImageNet and FC100 datasets, specifically in the 5-way 1-shot and 5-way 5-shot scenarios. Additionally, we performed cross-domain tasks across eight benchmark datasets, achieving high accuracy in the testing domains. These evaluations demonstrate the efficacy of our proposed method in comparison to state-of-the-art approaches. https://github.com/FatemehAskari/MSENet
@@ 1 Introduction
In recent years, deep learning has led to remarkable progress in the field of image recognition, significantly surpassing traditional computer vision algorithms bib1 ; fateh2024advancing . However, the success of deep learning models heavily relies on the availability of large datasets. When the available data is insufficient, models struggle to optimize their parameters effectively, leading to overfitting and ultimately hindering performance bib2 . This challenge is particularly pronounced in contexts where data labeling is time-consuming and costly, such as in medical imaging or rare object classification tian2024survey ; sun2024klsanet ; rezvani2024fusionlungnet . Therefore, the development of models capable of achieving acceptable performance with limited samples is critical bib7 .
@@
Data augmentation is one technique employed to mitigate the impact of limited labeled data wang2023data . However, traditional augmentation methods, such as rotation and noise addition, often fail to provide substantial new information, thereby limiting their effectiveness in preventing overfitting bib9 . Another approach, transfer learning bib10 , involves transferring knowledge from a source domain to a target domain by freezing shallow network layers while fine-tuning deeper layers. Yet, this method may struggle when the target domain significantly differs from the source bib11 .
To address these limitations, meta-learning has emerged as a promising solution li2023novel ; fateh2024msdnet . By leveraging prior learning experiences, meta-learning models can generalize across diverse tasks and rapidly adapt to new problem domains yang2024meta ; bib14 . The primary approaches in meta-learning include model-based, optimization-based, and metric-based methods bib15 . While model-based methods focus on architecture adjustments, optimization-based methods enhance learning through episodic training bib17 ; bib18 . Metric-based methods, however, learn a distance metric to measure sample similarity, ensuring that samples from the same class exhibit small distances liu2024few ; bib20 .
Despite their benefits, metric-based approaches typically rely on a single embedding space, which limits their ability to leverage the rich information from different feature representations. Recent advancements, such as the multi-distance metric network proposed by Gao et al. bib21 , suggest that utilizing multiple embedding spaces can enhance model performance by capturing both global and abstract features. Furthermore, the utilization of attention mechanisms and Transformers has significantly increased in recent years. For instance, methods like the SetFeat extractor introduced by Afrasiyabi et al. bib22 emphasize the importance of rich feature representations through self-attention mechanisms. These advancements highlight the growing recognition of the capabilities of attention mechanisms in enhancing accuracy and efficiency across various machine learning and computer vision tasks.
The ability to achieve high performance with limited labeled data highlights the practical value of our proposed approach, especially in scenarios where data collection and labeling are expensive or challenging. For instance, in the medical field, our model can assist in diagnosing rare diseases with only a small number of annotated samples. In industrial applications, it can support tasks such as anomaly detection or rare object classification in manufacturing lines. Similarly, in agriculture, it can facilitate the classification of plant species or pest detection with minimal labeled data.
@@
This research aims to propose a novel few-shot classification model that integrates various innovative components to enhance performance, particularly in scenarios where labeled data is scarce. Our model employs ResNet18 as a feature extractor, extracting feature maps from multiple stages to facilitate multi-scale representation. We introduce learnable parameter weights at each stage and incorporate self-attention mechanisms to enrich the feature space. Through comprehensive evaluations on the MiniImageNet and FC100 datasets, we demonstrate the effectiveness of our approach.
Our contributions can be summarized as follows:
@@
    We extract five feature maps from the backbone to capture both global and task-specific features.
@@
    We employ a self-attention mechanism for each feature map to capture more valuable information.
@@
    We incorporate learnable weights at each stage to enhance the model’s flexibility.
@@
    We propose a novel few-shot classification technique that significantly improves accuracy on the MiniImageNet and FC100 datasets.
@@ 2 Related Works
Koch et al. bib29 proposed a Siamese network that utilizes the VGG network as an extractor. They feed two pairs of images into the shared-weight convolutional network, and the network outputs a numerical value between 0 and 1, representing the similarity between the two images. Vinyals et al. bib24 proposed a matching network that computes the probability distribution over labels using an attention kernel. The attention kernel calculates the cosine similarity between the support set of embedded vectors and the query. It then normalizes the similarity using the softmax formula. Snell et al. bib30 proposed the Prototypical network, where each class in the support set is represented by a prototype, defined as the mean of the embedded vectors belonging to that class. The similarity between the query image’s embedded vector and the prototypes of each class is determined using the Euclidean distance. This enables the classification of query images into their respective classes. Sung et al. bib31 proposed the Relation Network, which does not rely on a separate distance function. Instead, it connects the representations of the support set and the query directly within the neural network architecture, allowing the network to learn the similarity measure. Previous few-shot image classification methods commonly used four-layer convolutional networks as backbones. However utilization of pre-trained networks such as ResNet12 and ResNet18 has become much more popular nowdays. However, calculating similarities and differences to a single feature vector is not sufficient. Gao et al. bib21 proposed a model called MDM-Net for few-shot learning. The MDM-Net maps input samples into four different feature spaces using a multi-output embedding network. Additionally, they introduced a task-adaptive margin to adjust the distance between different sample pairs. Transformers and attention mechanisms have emerged as state-of-the-art solutions in few-shot image classification, surpassing traditional CNN-based approaches. While CNNs have served as reliable feature extractors, their limitations, such as a restricted receptive field and parameter inefficiency, make them less effective in capturing complex patterns. In contrast, Transformers excel by capturing long-range dependencies, modeling non-local relationships, and efficiently parallelizing computations. Additionally, they offer enhanced interpretability by identifying key regions or features in the input data, providing insights into the model’s decision-making process. Recent advancements, such as the introduction of MDM-Net, demonstrate the superiority of Transformers and attention mechanisms in few-shot learning, combining their strengths to address the unique challenges of limited-data scenarios.
@@
Wang et al. bib32 propose a unified Query-Support Transformer (QSFormer) model for few-shot learning. The QSFormer model addresses the challenges of consistent image representations in both support and query sets, as well as effective metric learning between these sets. It consists of a sampleFormer branch that captures sample relationships and conducts metric learning using Transformer encoders, decoders, and cross-attention mechanisms. Additionally, a local patch Transformer (patchFormer) module is incorporated to extract structural representations from local image patches. The proposed model also introduces a Cross-scale Interactive Feature Extractor (CIFE) as an effective backbone module for extracting and fusing multi-scale CNN features. The QSFormer model demonstrates superior performance compared to existing methods in few-shot learning. Ran et al. bib32 propose a novel deep transformer and few-shot learning (DT-FSL) framework for hyperspectral image classification. The framework aims to achieve fine-grained classification using only a few-shot instances. By incorporating spatial attention and spectral query modules, the framework captures the relationships between non-local spatial samples and reduces class uncertainty. The network is trained using episodes and task-based learning strategies to enhance its modeling capability. Additionally, domain adaptation techniques are employed to reduce inter-domain distribution variation and achieve distribution alignment. Cheng et al. bib79 proposed a Class-Aware Patch Embedding Adaptation (CPEA) method for few-shot image classification, leveraging Vision Transformers (ViTs) pre-trained with Masked Image Modeling to generate semantically meaningful patch embeddings. They introduced class-aware embeddings to adapt patch embeddings, enabling class-relevant comparisons without explicit localization or alignment mechanisms, achieving state-of-the-art performance.
@@ 3.1 Problem definition
Figure 1: The overview architecture of the proposed model
The goal of few-shot classification is to classify a unseen sample. We have two datasets, $D_{train}$ and $D_{test}$, each associated with corresponding class sets $C_{train}$ and $C_{test}$, respectively. It is important the class sets $C_{train}$ and $C_{test}$ are disjoint, meaning they have no elements in common. Formally, we can express this as $C_{train}\cap C_{test}=\emptyset$. Each training episode consists of a support set $S$ and a query set $Q$. The support set $S$ comprises $K$ examples for each of $N$ distinct classes, denoted as $X_{i}^{S}=\left\{\left(x_{ij}^{s},C=i\right)\right\}_{i=1}^{K}$, where $x_{ij}^{s}$ represents the $j$th example belonging to class $i$. The query set $Q$ contains $X^{q}=\left\{x_{i}^{q}\right\}_{i=1}^{n_{q}}$. The objective of the model is to leverage the support set $S$ to correctly classify the query example $x_{i}^{q}$. In other words, the model is trained to predict the class label of the query image based on the set of supporting examples and their class affiliations provided in $S$. During training, each episode consists of a random sample drawn from the training dataset $D_{train}$. The objective of the model is to learn to extract feature representations from the examples in the support set $S$, such that the distance between the feature vector of the query image $x_{i}^{q}$ and the feature vectors of the support examples $x_{ij}^{s}$ can be effectively measured. Specifically, the model learns to extract discriminative feature vectors from the support examples, which can then be used to classify the query image based on its proximity to the support set features. This training paradigm encourages the model to rapidly adapt its feature extraction and classification capabilities from the limited support data to accurately predict the class label of the query instance. Similarly, during evaluation, the performance of the trained model is assessed on the held-out test dataset $D_{test}$. In this phase, the model extracts feature vectors for each example in the test set, leveraging the knowledge and feature extraction capabilities it learned during the training episodes on the $D_{train}$ dataset.
@@ 3.2 Overview
The overall architecture of our model is illustrated in Figure 1. Our proposed approach incorporates several key components designed to enhance performance. At the core is a robust backbone architecture that enables the extraction of feature maps across diverse spatial scales. Additionally, we have integrated an attention module to further refine the feature extraction process. Underpinning our framework is a distance metric that facilitates effective similarity computation between inputs. Moreover, we have incorporated learnable weights to capture the relative significance of each extracted feature map.
Figure 2: Visualization of feature maps from the five convolutional stages of ResNet-18, illustrating the progression from low-level features in shallow layers to high-level semantic features in deeper layers, critical for accurate classification.
@@ 3.2.1 Backbone
We utilized a pre-trained ResNet-18 network with initial weights from the ImageNet dataset. We removed the last fully-connected layer and fine-tuned the network on our specific dataset. To leverage multi-level feature maps, we proposed a multi-output embedding approach where we extracted feature maps at the end of each of the 5 convolutional blocks in the ResNet-18 architecture. This allowed us to capture feature representations at multiple scales and resolutions. As illustrated in Figure 2, the deeper layers of the ResNet-18 architecture play a more significant role in classification tasks compared to the shallow layers. While shallow layers capture low-level features such as edges and textures, the deeper layers focus on abstract and high-level semantic features that are crucial for distinguishing between classes. In our approach, we utilized the feature maps from all five stages of the ResNet-18 architecture to capture multi-scale representations. However, the deeper layers contribute more prominently to the final classification, as they extract the high-level semantic information critical for accurate class differentiation. Each sample in the support set, denoted as $x_{ij}^{s}$, is mapped to five different feature spaces like $f_{ij}^{s}$ as shown in Equation 1:
@@
In our proposed approach, we employ multiple convolutional layers $(Conv1,Conv2,Conv3,Conv4,Conv5)$ of the ResNet-18 network, which act as feature extractors to transform raw input images into meaningful and structured feature representations. A feature extractor refers to a mechanism in a neural network that automatically identifies and extracts important patterns or attributes (e.g., edges, textures, shapes, or semantic structures) from raw data. By leveraging these hierarchical feature maps from both the support and query samples, our method generates a robust, multi-level representation of the input images. These extracted features provide a compact and discriminative description of the images, facilitating effective similarity computation and accurate classification, which are crucial for our task.
Figure 3: SA module
@@ 3.2.2 Attention module
After extracting the feature vectors at each stage, we utilize self-attention and global average pooling. The representation of the attention module is illustrated in Figure 3 Considering the extracted feature vectors, we apply a $1\times 1$ convolution on $f$, resulting in convolutional vectors $k$, $g$ and $h$. This operation is performed to reduce the number of channels (Equation 3):
@@
The attention mechanism computes weights $\beta$, that determine the relative importance of each pixel in the feature map. These weights are calculated across the entire spatial extent, allowing the attention module to capture long-range dependencies beyond a local neighborhood, unlike traditional convolutions. The scaling factor, typically denoted as $\gamma$, is a learnable parameter in the network that is multiplied with the input feature $f^{Conv-p}$ before the addition (Equation 5):
$y^{Conv-p}=\gamma\times\beta_{i}+f^{Conv-p}$ (5)